OpenStudy (cakt07):

3 = 9 - --- 10 k + 12

OpenStudy (gorv):

\[\frac{ 3 }{ 10}=\frac{ 9 }{ k+12 }\]

OpenStudy (gorv):

step1: divide both side by 3

OpenStudy (gorv):

tell me what u got

OpenStudy (cakt07):

Like 3 / 3 over 10?

OpenStudy (gorv):

yeah

OpenStudy (cakt07):

im confused... .-. When i divide by both sides do you mean \[3\div 3/10\]

OpenStudy (cakt07):

so 3 div 3 is 1/10

OpenStudy (gorv):

yeah similarly divide on other side by 3

OpenStudy (cakt07):

3 div 9 is 3 over k + 12

OpenStudy (gorv):

yeah

OpenStudy (cakt07):

okay got it \[1/10 = 3/K + 12\]

OpenStudy (cakt07):

then what?

OpenStudy (gorv):

yeah now multiply both side by 10

OpenStudy (cakt07):

How do you know which number to divide and multiply though?

OpenStudy (cakt07):

\[1/10 \times 10 is 10 \over 10. \] and that cancels out also \[3/ K + 12 \times 10 = 10/ k + 12\]

OpenStudy (gorv):

we have to bring it to simple form

OpenStudy (cakt07):

im sorry 30/k + 12

OpenStudy (gorv):

\[\frac{ 1 }{ 10 }*10=\frac{ 3 }{ k+12 }*10\]

OpenStudy (gorv):

yeah

OpenStudy (cakt07):

Lol yeah

OpenStudy (cakt07):

and the one whole cancels out right?

OpenStudy (gorv):

yeah on left 10 cancel 10

OpenStudy (cakt07):

and now just solve 30/ K + 12

OpenStudy (cakt07):

do we divide\[30 \div 12?\]

OpenStudy (cakt07):

and the answer is K?

OpenStudy (gorv):

no no

OpenStudy (gorv):

we have now \[1=\frac{ 30 }{ k+12 }\]

OpenStudy (gorv):

multiply by k+12 both side

OpenStudy (cakt07):

so 12k x 30

OpenStudy (cakt07):

and 12k x 1?

OpenStudy (cakt07):

@gorv

OpenStudy (gorv):

k+12

OpenStudy (gorv):

\[k+12=\frac{ 30 }{ k+12 }*k+12\]